x^2-6x-27
_divided by_
x+3
When the teacher does it in class it looks so easy.I tried synthetic division but it does not come out right.I tried dividing -3 into x^2-6x-27 but I can not figure out how many times -3 will go into x which would actually be 1x,so I am stuck at the first step.

Answer by Theo(13342)

x^2 - 6x - 27 divided by x + 3

divide x into x^2 to get x ***** first part of your answer.

multiply x * (x+3) to get x^2 + 3x

subtract x^2 + 3x from x^2 - 6x to get -9x

carry down the -27 to get -9x - 27

divide x into -9x to get -9 ***** second part of your answer.

multiply -9 * (x+3) to get -9x - 27

subtract -9x - 27 from -9x - 27 to get 0.

your answer is:

x^2 - 6x - 27 divided by x + 3 = x - 9

This means that:

(x+3) * (x-9) = x^2 - 6x - 27

multiplying those factors out gets:

x^2 - 9x + 3x - 27 which becomes x^2 - 6x - 27 which is correct.

your answer is x - 9